Index

!, BASE-27 #[ ], BASE-30 <>, BASE-11 *, BASE-13 +, BASE-13 +, BASE-23 –, BASE-13 -, BASE-23 -D, TOOL-10 -batch, TOOL-9 -debug0, TOOL-10 -g, TOOL-10 -gcc, TOOL-9 -h, TOOL-9 -i, TOOL-9 -m, TOOL-9 -mo, TOOL-10 -noclean, TOOL-10 -norun, TOOL-10 -noshare, TOOL-9 -o, TOOL-9 -pervshare, TOOL-9 -r, TOOL-9 -run, TOOL-9 -runtime, TOOL-10 -target, TOOL-9 /, BASE-23 ::, BASE-18 :=, BASE-27 <, BASE-13 <, BASE-23 =, BASE-11 >, BASE-28 >, BASE-13 @, BASE-18 ^, BASE-29 ~, BASE-23

A EXEC SysIO, SYS-36 UnixPath, LIB-56

A READ SysIO, SYS-36 UnixPath, LIB-56

A WRITE SysIO, SYS-36 UnixPath, LIB-56

Abs, BASE-13

abs Integer, BASE-13 Real, BASE-23

access SysIO, SYS-36 UnixPath, LIB-56

add BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

add time, SYS-42

addCleaner

CleanUp, SYS-13 addition

Integer, BASE-13 Real, BASE-23

addToEnv

UnixEnv, LIB-55 AIX, TOOL-8 allocprof (CG), SYS-7 allocProfPrint

Control, SYS-18 allocProfReset

Control, SYS-18 alphac (CG), SYS-7 andb

Bits, BASE-4 ANTIQUOTE, BASE-20 app

BinaryDict, LIB-5 BinarySet, LIB-6 ByteArray, BASE-6 DICT, LIB-10 Fifo, LIB-14 IntMap, LIB-23 IntSet, LIB-24 List, BASE-18 ORD SET, LIB-36 Queue, LIB-43 SplayDict, LIB-46

SplaySet, LIB-47 append, BASE-15, GUIDE-10 appendPath

Pathname, LIB-38 apply

Code, SYS-14 HashTable, LIB-21 Name, LIB-33 PolyHashTable, LIB-41

architecture

System, SYS-40 arctan, BASE-23 argrep (CG), SYS-7 argv

CInterface, SYS-11 System, SYS-40

arithmetic, BASE-13 arithopt (CG), SYS-7 ARRAY, BASE-3 Array, BASE-3 array

Array, BASE-3 Array2, LIB-2 ByteArray, BASE-6 dynamic, LIB-12 DynamicArray, LIB-12 RealArray, BASE-26

ARRAY2, LIB-2 Array2, LIB-2 array2, LIB-2 ARRAY SORT, LIB-4 arrayoflist

Array, BASE-3 DynamicArray, LIB-12

ArrayQSort, LIB-4 associative arrays, LIB-10 associativity, GUIDE-10 AST, SYS-3 Ast, SYS-3 atime

SysIO, SYS-36 atob

StringCvt, LIB-49 atof

StringCvt, LIB-49 atoi

StringCvt, LIB-49 BadArg, LIB-1 badArg, LIB-1 BadArgList

Format, LIB-17 BadFormat

Format, LIB-17 BadInput

Format, LIB-17 BadPrecision

Makestring, LIB-31 base definitions, LIB-1 baseOfPath

Pathname, LIB-38 batoi

StringCvt, LIB-49 before, BASE-11 betacontract (CG), SYS-7 betaexpand (CG), SYS-7 Bin

StringCvt, LIB-49 binary tree, LIB-6 BinaryDict, LIB-5 BinarySet, LIB-6 Bind, BASE-11 bindContainsVar

MC, SYS-23 bindExhaustive

MC, SYS-23 BITS, BASE-4 Bits, BASE-4 bitwise operations, BASE-4 blas

CInterface, SYS-11 blast read

Unsafe, SYS-44 blast write

Unsafe, SYS-44 bodysize (CG), SYS-7 BOOL, BASE-5

Format, LIB-17 Bool, BASE-5 bool

Bool, BASE-5 Boolean, BASE-5 boolToStr

Makestring, LIB-31 boot

Unsafe, SYS-44 bound

DynamicArray, LIB-12 boxed

Unsafe, SYS-44 Boxity

Unsafe, SYS-44 branchfold (CG), SYS-7 bsd, TOOL-9 bsd386, TOOL-8 bstore

Unsafe, SYS-44 bucketSizes

HashTable, LIB-21 PolyHashTable, LIB-41

bugs, GUIDE-1 BYTEARRAY, BASE-6 ByteArray, BASE-6 bytearray

ByteArray, BASE-6

c function

CInterface, SYS-11 c string

CInterface, SYS-11 call with current continuation, BASE-

8

callcc, BASE-8

PolyCont, SYS-24 calleeGen (CG), SYS-8 calleesaves (CG), SYS-8 can input, BASE-15 capture

PolyCont, SYS-24 cast

Unsafe, SYS-44 catalogEnv

Environment, SYS-21 cd

Directory, SYS-20 ceiling, BASE-23 CFunNotFound

CInterface, SYS-11 CG, SYS-7

Control, SYS-18 CGCONTROL, SYS-7 changeLvars

Compile, SYS-15 CHAR SET, LIB-7 char set

CharSet, LIB-7 character classification, LIB-8 character sets, LIB-7 character strings, BASE-28 CharSet, LIB-7 charSetOfList

CharSet, LIB-7 charsOfSet

CharSet, LIB-7 chdir

CInterface, SYS-11 check timer, SYS-42 check timer gc, SYS-42 check timer sys, SYS-42 chmod

SysIO, SYS-36 Chr

String, BASE-28 chr

String, BASE-28 CINTERFACE, SYS-11 CInterface, SYS-11 clean mode

CleanUp, SYS-13 CleanForExportFn

CleanUp, SYS-13 CleanForExportML

CleanUp, SYS-13 CleanForInit

CleanUp, SYS-13 CleanForQuit

CleanUp, SYS-13 CLEANUP, SYS-13 CleanUp, SYS-13 cleanup

CleanUp, SYS-13 close

SysIO, SYS-36 close in, BASE-15 close out, BASE-15 closeSource

Compile, SYS-15 closureprint (CG), SYS-7 closureStrategy (CG), SYS-7 CML, TOOL-3 cmpKey

ORD KEY, LIB-35 SplayDict, LIB-46 SplaySet, LIB-47

CODE, SYS-14 code, SYS-14

Code, SYS-14 code generator, SYS-7 code-generator generator, TOOL-12 codeUnit

Compile, SYS-15 collected

Runtime, SYS-31 collectedfrom

Runtime, SYS-31 column

Array2, LIB-2 comment (CG), SYS-7 comparefold (CG), SYS-8 compilation unit, GUIDE-10 compilation, separate, TOOL-16 COMPILE, SYS-15 compile

Compile, SYS-15 compileAst

Compile, SYS-15 compiler interface, SYS-15 compressStr

StringUtil, LIB-51 compUnit

Compile, SYS-15 concatEnv

Environment, SYS-21 Concurrent ML, TOOL-3 connect inet

SysIO, SYS-36 connect unix

SysIO, SYS-36 cont, BASE-8

PolyCont, SYS-24 contents

Fifo, LIB-14 Queue, LIB-43

continuation, BASE-8, SYS-32 CONTROL, SYS-18 Control, SYS-18 control cont

PolyCont, SYS-24 copy

HashTable, LIB-21 Name, LIB-33 PolyHashTable, LIB-41

cos, BASE-23

cpsopt (CG), SYS-7

create b

Unsafe, SYS-44 create s

Unsafe, SYS-44 cross compiler, TOOL-9 cse (CG), SYS-8 csehoist (CG), SYS-8 cspan

StringUtil, LIB-51 cspanp

StringUtil, LIB-51 ctime

SysIO, SYS-36 CTYPE, LIB-8 CType, LIB-8

datalist

Unsafe, SYS-44 deadvars (CG), SYS-7 Debug

Control, SYS-18 debug, TOOL-9 debugBind

Control, SYS-18 debugCollect

Control, SYS-18 debugger, TOOL-4 debugging

Control, SYS-18 debugLook

Control, SYS-18 Dec

StringCvt, LIB-49 dec

Ref, BASE-27 default

DynamicArray, LIB-12 defaultExt

Pathname, LIB-38 Definition of Standard ML, GUIDE-9 delete BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

Dequeue Fifo, LIB-14 Queue, LIB-43

dequeue Fifo, LIB-14 Queue, LIB-43

DESC

SysIO, SYS-36 describe

Environment, SYS-21 DICT, LIB-10 dict

BinaryDict, LIB-5 DICT, LIB-10 SplayDict, LIB-46

dictionaries

applicative, LIB-10 dictionary

splay tree, LIB-46 Diff, BASE-13, GUIDE-10 difference

BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

dimensions

Array2, LIB-2 DIRECTORY, SYS-20 Directory, SYS-20 dirOfPath

Pathname, LIB-38 Div, BASE-13, GUIDE-10 div, BASE-13 dropargs (CG), SYS-7 dtablesize

SysIO, SYS-36 dup

SysIO, SYS-36 DYNAMIC ARRAY, LIB-12 DynamicArray, LIB-12 dynix, TOOL-8

earlier, SYS-42 elaborate

Compile, SYS-15 Emacs, GUIDE-4 empty

BinarySet, LIB-6 Fifo, LIB-14 IntMap, LIB-23 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

emptyEnv

Environment, SYS-21 end of stream, BASE-15 enqueue

Fifo, LIB-14 Queue, LIB-43

Env, SYS-21 environ

CInterface, SYS-11 System, SYS-40 UnixEnv, LIB-55

ENVIRONMENT, SYS-21 environment

Environment, SYS-21 name-value (Unix), LIB-55

EOF, GUIDE-2 Equal, LIB-1 escape

PolyCont, SYS-24 escapeGen (CG), SYS-8 eta (CG), SYS-7 etasplit (CG), SYS-8 eval stream

Compile, SYS-15 exceptions, polymorphic, GUIDE-6 exec

CInterface, SYS-11 execute

Compile, SYS-15 IO, BASE-15

execute in env

IO, BASE-15 execution profiler, TOOL-14 execution timing, SYS-42 eXene, TOOL-5 exists

List, BASE-18 exit

CInterface, SYS-11 exn, BASE-11 exn name

System, SYS-40 Exp, BASE-23, GUIDE-10 exp, BASE-23 expandResult

MC, SYS-23 expandStr

StringUtil, LIB-51 explode

String, BASE-28 export

CInterface, SYS-11 exportFn, BASE-10 exportML, BASE-10 EXPRESSIONS, SYS-3 extOfPath

Pathname, LIB-38 extract

ByteArray, BASE-6 extraflatten (CG), SYS-8

F BLK SysIO, SYS-36 UnixPath, LIB-56

F CHR SysIO, SYS-36 UnixPath, LIB-56

F DIR SysIO, SYS-36 UnixPath, LIB-56

F REGULAR SysIO, SYS-36 UnixPath, LIB-56

F SOCK SysIO, SYS-36 UnixPath, LIB-56

F SYMLINK SysIO, SYS-36 UnixPath, LIB-56

Fail, BASE-11 false

Bool, BASE-5 fastlib, GUIDE-11 FCTspace, SYS-34 fctSymbol, SYS-34 fd

SysIO, SYS-36 FIFO, LIB-14 Fifo, LIB-14 fifo

Fifo, LIB-14 file, BASE-10 file names, LIB-38 file type

SysIO, SYS-36 UnixPath, LIB-56

fileid

SysIO, SYS-36 files, SYS-20 filter

HashTable, LIB-21 ListUtil, LIB-29 Name, LIB-33 PolyHashTable, LIB-41

filterEnv

Environment, SYS-21 filterStaticEnv

Environment, SYS-21 finalize

Finalizer, LIB-16 FINALIZER, LIB-16 Finalizer, LIB-16 find

BinaryDict, LIB-5 BinarySet, LIB-6 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 ListUtil, LIB-29 Name, LIB-33 ORD SET, LIB-36 PolyHashTable, LIB-41 SplayDict, LIB-46 SplaySet, LIB-47

findFile Pathname, LIB-38 UnixPath, LIB-56

findFileOfType

UnixPath, LIB-56 findFiles

Pathname, LIB-38 UnixPath, LIB-56

findFilesOfType

UnixPath, LIB-56 findOne

ListUtil, LIB-29 findstr

StringUtil, LIB-51 fionread

SysIO, SYS-36 fixed precision, BASE-13 FIXspace, SYS-34 fixSymbol, SYS-34 flatten

ListUtil, LIB-29 flattenargs (CG), SYS-7 floatargs (CG), SYS-8 floating-point, BASE-23 floatreg (CG), SYS-8 floatvars (CG), SYS-8 Floor, BASE-23, GUIDE-10 floor, BASE-23 flush

Print, SYS-29 flush out, BASE-15 fmt item

Format, LIB-17 fname

SysIO, SYS-36 fold

BinaryDict, LIB-5 BinarySet, LIB-6 ByteArray, BASE-6 DICT, LIB-10 Fifo, LIB-14 IntMap, LIB-23 IntSet, LIB-24 List, BASE-18 ORD SET, LIB-36 Queue, LIB-43 SplayDict, LIB-46 SplaySet, LIB-47

foldconst (CG), SYS-8 FORMAT, LIB-17 Format, LIB-17 format

Format, LIB-17 formatf

Format, LIB-17 formatList

ListFormat, LIB-26 formatting lists, LIB-26 frag

Quote, BASE-20 fromTo

ListUtil, LIB-29 FSIGspace, SYS-34 fsigSymbol, SYS-34 fsize

SysIO, SYS-36 ftype

SysIO, SYS-36

garbage collector, SYS-31 gc

CInterface, SYS-11 gcc, TOOL-9 gcmessages

Runtime, SYS-31 GENERAL, BASE-11 General, BASE-8 genList

ListUtil, LIB-29 getDead

Finalizer, LIB-16 getdirent

SysIO, SYS-36 getEnv

UnixEnv, LIB-55 getfid

SysIO, SYS-36 getFromEnv

UnixEnv, LIB-55 getgid

CInterface, SYS-11 gethdlr

Unsafe, SYS-44 gethostid

CInterface, SYS-11 gethostname

CInterface, SYS-11 getmod

SysIO, SYS-36 getObjTag

Unsafe, SYS-44 getownid

SysIO, SYS-36 getPath

UnixPath, LIB-56 getpid

CInterface, SYS-11 getSpecial

Unsafe, SYS-44 gettime

CInterface, SYS-11 gettimeofday

CInterface, SYS-11 getuid

CInterface, SYS-11 getValue

UnixEnv, LIB-55 getvar

Unsafe, SYS-44 getWD

Directory, SYS-20 UnixPath, LIB-56

GNU C, TOOL-9 Greater, LIB-1 greater

Integer, BASE-13 Real, BASE-23 String, BASE-28

handlerfold (CG), SYS-7 hash

strings, LIB-20 table, LIB-21 tables, polymorphic, LIB-41

HASH KEY, LIB-21 hash key

HashTable, LIB-21 HASH TABLE, LIB-21 hash table

HashTable, LIB-21 PolyHashTable, LIB-41

HashString, LIB-20 hashString, LIB-20 HashTable, LIB-21 hashVal

HashTable, LIB-21 Hd

List, BASE-18 hd

List, BASE-18 head

Fifo, LIB-14

Queue, LIB-43 heap resizing, SYS-31 heapsize, TOOL-9 hoist (CG), SYS-7 hoistdown (CG), SYS-7 hoistup (CG), SYS-7 HPPA, TOOL-8 HPUX, TOOL-8

i386, TOOL-8 i486, TOOL-8 icount (CG), SYS-8 ifidiom (CG), SYS-8 implode

String, BASE-28 Impossible, LIB-1 inc

Ref, BASE-27 index

StringCvt, LIB-49 StringUtil, LIB-51

indexable, BASE-3 indexable sequences, BASE-30 indexing

Control, SYS-18 indexp

StringUtil, LIB-51 INFO, TOOL-6 Info, TOOL-6 info, TOOL-6 input, BASE-15 Input/Output, BASE-15, SYS-36 input line, BASE-15 inputc, BASE-15 inputCode

Code, SYS-14 inqHandler

Signals, SYS-32 insert

BinaryDict, LIB-5 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 Name, LIB-33 PolyHashTable, LIB-41 SplayDict, LIB-46

inSet CharSet, LIB-7 inSetOrd

CharSet, LIB-7 instream, BASE-15 instSigs

Control, SYS-18 INT

Format, LIB-17 int, BASE-13

Unsafe, SYS-44 INTEGER, BASE-13 Integer, BASE-13 integer map, LIB-23 integer set, LIB-24 Intel, TOOL-8 interactive

System, SYS-40 internals

Control, SYS-18 interp

Control, SYS-18 interpreter, TOOL-9 Interrupt, BASE-11 interrupt, GUIDE-2, SYS-32 intersection

BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

INTMAP, LIB-23 IntMap, LIB-23 intmap

IntMap, LIB-23 INTSET, LIB-24 IntSet, LIB-24 intset

IntSet, LIB-24 intToBin

Makestring, LIB-31 intToHex

Makestring, LIB-31 intToOct

Makestring, LIB-31 intToStr

Makestring, LIB-31 invariant (CG), SYS-7 IO, BASE-10 Io, BASE-15, GUIDE-10 irix, TOOL-8 is term in, BASE-15 is term out, BASE-15 isAbsolutePath

Pathname, LIB-38 isAlpha, LIB-8 isAlphaNum, LIB-8 isAlphaNumOrd, LIB-8 isAlphaOrd, LIB-8 isAscii, LIB-8 isAsciiOrd, LIB-8 isatty

SysIO, SYS-36 isCntrl, LIB-8 isCntrlOrd, LIB-8 isDigit, LIB-8 isDigitOrd, LIB-8 isDir

Directory, SYS-20 isEmpty

Fifo, LIB-14 Queue, LIB-43

isGraph, LIB-8 isGraphOrd, LIB-8 isLower, LIB-8 isLowerOrd, LIB-8 isPrefix

StringUtil, LIB-51 isPrint, LIB-8 isPrintOrd, LIB-8 isPunct, LIB-8 isPunctOrd, LIB-8 isSpace, LIB-8 isSpaceOrd, LIB-8 isUpper, LIB-8 isUpperOrd, LIB-8 isXDigit, LIB-8 isXDigitOrd, LIB-8 item

ORD SET, LIB-36 SplaySet, LIB-47

ITERATE, LIB-25 Iterate, LIB-25 iterate

Iterate, LIB-25 iteration, LIB-25

join

SplayTree, LIB-48

Key HashTable, LIB-21 SplayDict, LIB-46

keys, ordered, LIB-35 kind

Symbol, SYS-34 knownClGen (CG), SYS-8 knownfiddle (CG), SYS-7 knownGen (CG), SYS-8

L INCR

SysIO, SYS-36 L SET

SysIO, SYS-36 L XTND

SysIO, SYS-36 LABspace, SYS-34 labSymbol, SYS-34 lambdaopt (CG), SYS-7 lambdaprop (CG), SYS-7 lastratio

Runtime, SYS-31 layerEnv

Environment, SYS-21 layerStatic

Environment, SYS-21 LEFT

Format, LIB-17 length

Array, BASE-3 ByteArray, BASE-6 Fifo, LIB-14 List, BASE-18 Queue, LIB-43 RealArray, BASE-26 String, BASE-28 Vector, BASE-30

Less, LIB-1 less than

Integer, BASE-13 Real, BASE-23 String, BASE-28

Lex, TOOL-11 lexical analyzer, TOOL-11 LIB BASE, LIB-1 LibBase, LIB-1 line, BASE-15 linewidth

Print, SYS-29 link

SysIO, SYS-36 LIST, BASE-18 List, BASE-18 list, BASE-18

formatted, LIB-26

utility functions, LIB-29 LIST FORMAT, LIB-26 list UTIL, LIB-29 listDir

Directory, SYS-20 ListFormat, LIB-26 listItems

BinaryDict, LIB-5 BinarySet, LIB-6 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 IntSet, LIB-24 Name, LIB-33 ORD SET, LIB-36 PolyHashTable, LIB-41 SplayDict, LIB-46 SplaySet, LIB-47

ListMergeSort, LIB-28 ListUtil, LIB-29 Ln, BASE-23 ln, BASE-23 loading source files, GUIDE-2 local, GUIDE-10 logical operations, BASE-4 lookahead, BASE-15 lookup

Unsafe, SYS-44 lseek

SysIO, SYS-36 lshift

Bits, BASE-4

mach, TOOL-8 machdep (CG), SYS-7 machine code, SYS-14 mailing list, GUIDE-1 majorcollections

Runtime, SYS-31 make, TOOL-16 makeml, TOOL-8 makePath

Pathname, LIB-38 makeSource

Compile, SYS-15 MAKESTRING, LIB-31 Makestring, LIB-31 makestring, SYS-42

Bool, BASE-5 Integer, BASE-13 Real, BASE-23 Symbol, SYS-34

makeSymbol

Symbol, SYS-34 map

BinaryDict, LIB-5 DICT, LIB-10 Fifo, LIB-14 HashTable, LIB-21 IntMap, LIB-23 List, BASE-18 Name, LIB-33 PolyHashTable, LIB-41 Queue, LIB-43 SplayDict, LIB-46

markabsyn

Control, SYS-18 masked

Signals, SYS-32 maskSignals

Signals, SYS-32 Match, BASE-11 matchExhaustive

MC, SYS-23 matchRedundant

MC, SYS-23 max

Integer, BASE-13 maxregs (CG), SYS-7 MC, SYS-23

Control, SYS-18 MC680x0, TOOL-8 MCCONTROL, SYS-23 member

BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

merge sort, LIB-28 messages

garbage collection, SYS-31 min

Integer, BASE-13 minorcollections

Runtime, SYS-31 minus

Integer, BASE-13 Real, BASE-23

Mips, TOOL-8 misc1 (CG), SYS-7 misc2 (CG), SYS-7 misc3 (CG), SYS-7 misc4 (CG), SYS-7 mkAbsolutePath

Pathname, LIB-38 UnixPath, LIB-56

mkCharSet

CharSet, LIB-7 mkCode

Code, SYS-14 mkDict

BinaryDict, LIB-5 DICT, LIB-10 SplayDict, LIB-46

mkdir

SysIO, SYS-36 mkName

Name, LIB-33 mkNameTbl

Name, LIB-33 mkQueue

Queue, LIB-43 mkRandom

Random, LIB-45 mkRelativePath

Pathname, LIB-38 UnixPath, LIB-56

mkSearchPath

Pathname, LIB-38 mkTable

HashTable, LIB-21 PolyHashTable, LIB-41

ML-Lex, TOOL-11 ML-Twig, TOOL-12 ML-Yacc, TOOL-13 Mod, BASE-13, GUIDE-10 mod, BASE-13 mode

SysIO, SYS-36 module, higher-order, GUIDE-6 mtime

SysIO, SYS-36 multiplication

Integer, BASE-13 Real, BASE-23

NAME, LIB-33 Name, LIB-33 name

Name, LIB-33 Symbol, SYS-34

name tbl

Name, LIB-33 nameOfPath

Pathname, LIB-38 nameSpace

Symbol, SYS-34 namespace

Symbol, SYS-34 Neg, BASE-13, GUIDE-10 negate

Integer, BASE-13 Real, BASE-23

netnews, GUIDE-1 newconreps (CG), SYS-7 NeXT, TOOL-8 NeXTstep, TOOL-8 nil, BASE-18 NONE

Bool, BASE-5 norm

Random, LIB-45 noshare, BASE-10 NoSuchFile

Pathname, LIB-38 not

Bool, BASE-5 notb

Bits, BASE-4 NotDirectory

Directory, SYS-20 NotFound

BinaryDict, LIB-5 DICT, LIB-10 IntMap, LIB-23 IntSet, LIB-24 ORD SET, LIB-36 SplayDict, LIB-46 SplaySet, LIB-47 StringUtil, LIB-51

Nth

List, BASE-18 nth

List, BASE-18 NthTail

List, BASE-18 nthtail

List, BASE-18 null

List, BASE-18 numItems

BinaryDict, LIB-5 BinarySet, LIB-6 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 IntSet, LIB-24 Name, LIB-33 ORD SET, LIB-36 PolyHashTable, LIB-41 SplayDict, LIB-46 SplaySet, LIB-47

o, BASE-11 O APPEND

SysIO, SYS-36 O READ

SysIO, SYS-36 O WRITE

SysIO, SYS-36 oatoi

StringCvt, LIB-49 Obj

Finalizer, LIB-16 object

Finalizer, LIB-16 objinfo

Finalizer, LIB-16 objLength

Unsafe, SYS-44 Oct

StringCvt, LIB-49 open, GUIDE-10 open append, BASE-15 open in, BASE-15 open out, BASE-15 open string, BASE-15 openf

SysIO, SYS-36 operating system interface, SYS-11 optafterclosure (CG), SYS-8 optimization, GUIDE-10 optimizer control, SYS-7 option

Bool, BASE-5 orb

Bits, BASE-4 Ord

String, BASE-28 ord

String, BASE-28 ORD KEY, LIB-35 ord key

ORD KEY, LIB-35 SplayDict, LIB-46 SplaySet, LIB-47

ORD SET, LIB-36 ordered keys, LIB-35 ordof

String, BASE-28 Unsafe, SYS-44

out

Print, SYS-29 output, BASE-15 outputc, BASE-15 outstream, BASE-15 Overflow, BASE-13, GUIDE-10

padLeft

Makestring, LIB-31 padRight

Makestring, LIB-31 parse

Compile, SYS-15 parser generator, TOOL-13 PATH

SysIO, SYS-36 path, SYS-3

Pathname, LIB-38 path (CG), SYS-7 path, search, LIB-56 path list

UnixPath, LIB-56 pathExplode

Pathname, LIB-38 pathImplode

Pathname, LIB-38 PATHNAME, LIB-38 Pathname, LIB-38 pathname decomposition, LIB-38 pathsOfList

Pathname, LIB-38 pattern-match compiler, SYS-23 pause

Signals, SYS-32 peek

BinaryDict, LIB-5 BinarySet, LIB-6 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 Name, LIB-33 ORD SET, LIB-36 PolyHashTable, LIB-41 SplayDict, LIB-46 SplaySet, LIB-47

pervasiveEnvRef

Environment, SYS-21 pipe

SysIO, SYS-36 plus

Integer, BASE-13 Real, BASE-23 POLY CONT, SYS-24 POLY HASH TABLE, LIB-41 PolyCont, SYS-24 PolyHashTable, LIB-41 polymorphic exceptions, GUIDE-6 pp (prettyprinting), SYS-25 ppconsumer, SYS-25 ppstream, SYS-25 prefix

ListUtil, LIB-29 prefixCmp

StringUtil, LIB-51 prependPath

Pathname, LIB-38 PRETTYPRINT, SYS-25 prettyprinting, SYS-25 primaryPrompt

Control, SYS-18 Print, SYS-29

Control, SYS-18 print

Bool, BASE-5 Integer, BASE-13 Real, BASE-23 String, BASE-28

printArgs

MC, SYS-23 PRINTCONTROL, SYS-29 printDepth

Print, SYS-29 printit (CG), SYS-8 printLambda (CG), SYS-8 printLength

Print, SYS-29 printLoop

Print, SYS-29 printRet

MC, SYS-23 printsize (CG), SYS-8 prLambda

Control, SYS-18 Prod, BASE-13, GUIDE-10 PROFILE, TOOL-14 Profile, TOOL-14 profileOff

Profile, TOOL-14 profileOn

Profile, TOOL-14 profiling

Profile, TOOL-14 prompt, GUIDE-2 pstruct

Unsafe, SYS-44

QUEUE, LIB-43 Queue, LIB-43 queue

applicative, LIB-14 imperative, LIB-43

Queue, LIB-43 quicksort, LIB-4 Quot, BASE-13 quot, BASE-13 quotation, BASE-20

Control, SYS-18 QUOTE, BASE-20

radix

StringCvt, LIB-49 RANDOM, LIB-45 Random, LIB-45 random

Random, LIB-45 random number, LIB-45 Range

ByteArray, BASE-6 range

Random, LIB-45 rangeopt (CG), SYS-8 ratio, TOOL-9

Runtime, SYS-31 read

SysIO, SYS-36 readi

SysIO, SYS-36 readlink

SysIO, SYS-36 REAL, BASE-23

Format, LIB-17 Real, BASE-23 real, BASE-23

Unsafe, SYS-44 RealArray, BASE-26 realarray, BASE-26 realEFormat

Makestring, LIB-31 realFFormat

Makestring, LIB-31 realGFormat

Makestring, LIB-31 RealSubscript

RealArray, BASE-26 realToFloStr

Makestring, LIB-31 realToSciStr

Makestring, LIB-31 recordcopy (CG), SYS-7 recordopt (CG), SYS-7 recordpath (CG), SYS-7 reduce (CG), SYS-7 reducemore (CG), SYS-7 REF, BASE-27 Ref, BASE-27 References, GUIDE-4 references, BASE-27 registerObj

Finalizer, LIB-16 relation, LIB-1 rem, BASE-13 remove

BinaryDict, LIB-5 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 ListUtil, LIB-29 Name, LIB-33 PolyHashTable, LIB-41 SplayDict, LIB-46

removeCleaner

CleanUp, SYS-13 removeFromEnv

UnixEnv, LIB-55 removeOne

ListUtil, LIB-29 reopen

Control, SYS-18 repeat

Iterate, LIB-25 report

Profile, TOOL-14 representations (CG), SYS-8 reset

Profile, TOOL-14 rev

List, BASE-18 revapp

BinaryDict, LIB-5 BinarySet, LIB-6 ByteArray, BASE-6 DICT, LIB-10 Fifo, LIB-14 IntMap, LIB-23 IntSet, LIB-24 List, BASE-18 ORD SET, LIB-36 Queue, LIB-43 SplayDict, LIB-46 SplaySet, LIB-47

revfindstr

StringUtil, LIB-51 revfold

BinaryDict, LIB-5 BinarySet, LIB-6 ByteArray, BASE-6 DICT, LIB-10 Fifo, LIB-14 IntMap, LIB-23 IntSet, LIB-24 List, BASE-18 ORD SET, LIB-36 Queue, LIB-43 SplayDict, LIB-46 SplaySet, LIB-47

revindex

StringUtil, LIB-51 revindexp

StringUtil, LIB-51 RIGHT

Format, LIB-17 rounds (CG), SYS-7 row

Array2, LIB-2 RS/6000, TOOL-8 rshift

Bits, BASE-4 Runtime, SYS-31

Control, SYS-18 runtime system, SYS-31 RUNTIMECONTROL, SYS-31 runtimeStamp

System, SYS-40

salb

CInterface, SYS-11 sameKey

HashTable, LIB-21 sameName

Name, LIB-33 saveLambda

Control, SYS-18 saveLvarNames

Control, SYS-18 say

Print, SYS-29 scan

Format, LIB-17 scani

Format, LIB-17 ScanList

ListFormat, LIB-26 scanList

ListFormat, LIB-26 scheduling (CG), SYS-8 search, TOOL-6 secondaryPrompt

Control, SYS-18 select

SysIO, SYS-36 selectopt (CG), SYS-7 send obd

SysIO, SYS-36 separate compilation, TOOL-16 sequences, BASE-3 sequent, TOOL-8 set

of characters, LIB-7 of integers, LIB-24 on ordered types, LIB-36 ORD SET, LIB-36 splay tree, LIB-47 SplaySet, LIB-47

set, binary tree, LIB-6 set term in, BASE-15 set term out, BASE-15 setHandler

Signals, SYS-32 sethdlr

Unsafe, SYS-44 setitimer

CInterface, SYS-11 setSpecial

Unsafe, SYS-44 setvar

Unsafe, SYS-44 SGI, TOOL-8 shutdown

CleanUp, SYS-13 SIGALRM, SYS-32 SIGCHLD, SYS-32 SIGCONT, SYS-32 SIGGC, SYS-32 SIGHUP, SYS-32 SIGINT, SYS-32 SIGIO, SYS-32 SIGNALS, SYS-32 Signals, SYS-32 signature, SYS-42 signatures

Print, SYS-29 SIGPROF, SYS-32 SIGQUIT, SYS-32 SIGspace, SYS-34 sigSymbol, SYS-34 SIGTERM, SYS-32 SIGTSTP, SYS-32 SIGURG, SYS-32 SIGUSR1, SYS-32 SIGUSR2, SYS-32 SIGVTALRM, SYS-32 SIGWINCH, SYS-32 sin, BASE-23 singleton

BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

Size Array, BASE-3 ByteArray, BASE-6 DynamicArray, LIB-12 RealArray, BASE-26 Vector, BASE-30

size

String, BASE-28 sizeOf

Code, SYS-14 slength

Unsafe, SYS-44 sml, GUIDE-2 smlc, TOOL-9 smld, TOOL-9 smli, TOOL-9 softlimit, TOOL-9 softmax

Runtime, SYS-31 SOME

Bool, BASE-5 sort, LIB-4

ARRAY SORT, LIB-3 ArrayQSort, LIB-4 LIST SORT, LIB-27 ListMergeSort, LIB-28

sorted, LIB-4 ARRAY SORT, LIB-3 ArrayQSort, LIB-4 LIST SORT, LIB-27 ListMergeSort, LIB-28

sorting, LIB-3 array, LIB-4 list, LIB-28

SourceGroup, GUIDE-3, TOOL-16 span

StringUtil, LIB-51 spanp

StringUtil, LIB-51 SPARC, TOOL-8 special

Unsafe, SYS-44 spillGen (CG), SYS-8 splay

SplayTree, LIB-48 splay tree, LIB-48 SPLAY TREE, LIB-48 SplayDict, LIB-46 SplayNil

SplayTree, LIB-48 SplayObj

SplayTree, LIB-48 SplaySet, LIB-47 SplayTree, LIB-48 Split

ListUtil, LIB-29 split

ListUtil, LIB-29 splitp

ListUtil, LIB-29 splitPath

Pathname, LIB-38 Sqrt, BASE-23 sqrt, BASE-23 STATIC ARRAY, LIB-12 staticEnv

Environment, SYS-21 staticPart

Environment, SYS-21 staticUnit

Compile, SYS-15 std err, BASE-15 std in, BASE-15 std out, BASE-15 store

Unsafe, SYS-44 store s

Unsafe, SYS-44 STR

Format, LIB-17 strcmp

StringUtil, LIB-51 STRING, BASE-28 String, BASE-28 string, BASE-15

conversion to primitive values, LIB-

49 Makestring, LIB-31 operations, LIB-51 StringUtil, LIB-51 unique, LIB-33 Unsafe, SYS-44

STRING CVT, LIB-49 STRING UTIL, LIB-51 StringCvt, LIB-49 stringDepth

Print, SYS-29 stringMap

StringUtil, LIB-51 stringOf

Name, LIB-33 strings

formatted, LIB-17 stringTrans

StringUtil, LIB-51 StringUtil, LIB-51 strong

Weak, SYS-47 STRspace, SYS-34 strSymbol, SYS-34 strToBool

StringCvt, LIB-49 strToInt

StringCvt, LIB-49 strToReal

StringCvt, LIB-49 sub

Array, BASE-3 Array2, LIB-2 ByteArray, BASE-6 DynamicArray, LIB-12 RealArray, BASE-26 Vector, BASE-30

sub time, SYS-42 subArray

DynamicArray, LIB-12 Subscript

Array, BASE-3 ByteArray, BASE-6 DynamicArray, LIB-12 RealArray, BASE-26 Vector, BASE-30

subscript

Unsafe, SYS-44 subscriptf

Unsafe, SYS-44 subscriptv

Unsafe, SYS-44 Substring

String, BASE-28 substring

String, BASE-28 subtraction

Integer, BASE-13 Real, BASE-23

suffix ListUtil, LIB-29 StringUtil, LIB-51

Sum, BASE-13, GUIDE-10 Sun 3, TOOL-8 sunos, TOOL-8 switchopt (CG), SYS-7 SYMBOL, SYS-34 Symbol, SYS-34 symlink

SysIO, SYS-36 syntax tree, SYS-3 syscall

CInterface, SYS-11 SysError

CInterface, SYS-11 SYSIO, SYS-36 SysIO, SYS-36 SYSTEM, SYS-40 System, SYS-40 system

CInterface, SYS-11 System, SYS-40

SystemCall CInterface, SYS-11

tables, LIB-10 tabulate

Array, BASE-3 Array2, LIB-2 DynamicArray, LIB-12 Vector, BASE-30

tagopt (CG), SYS-7 tail (CG), SYS-7 tailrecur (CG), SYS-7 targeting (CG), SYS-7 throw, BASE-8

PolyCont, SYS-24 TIME, SYS-42 time

CInterface, SYS-11 Timer, SYS-42

TimeLimit, LIB-54 timeLimit

TimeLimit, LIB-54 TimeOut

TimeLimit, LIB-54 Timer, SYS-42 timer, SYS-42 timings

Control, SYS-18 Tl

List, BASE-18 tl

List, BASE-18 toAscii, LIB-8 toAsciiOrd, LIB-8 tokenize

StringUtil, LIB-51 tokenizep

StringUtil, LIB-51 toLower, LIB-8 toLowerOrd, LIB-8 toplevelcont

Unsafe, SYS-44 topLevelEnvRef

Environment, SYS-21 toUpper, LIB-8 toUpperOrd, LIB-8 transform

BinaryDict, LIB-5 DICT, LIB-10 HashTable, LIB-21 IntMap, LIB-23 Name, LIB-33 PolyHashTable, LIB-41 SplayDict, LIB-46

trapv (CG), SYS-7 true

Bool, BASE-5 truncate, BASE-23

SysIO, SYS-36 tuple

Unsafe, SYS-44 Twig, TOOL-12 TYCspace, SYS-34 tycSymbol, SYS-34 TYVspace, SYS-34 tyvSymbol, SYS-34

ultrix, TOOL-8 umask

SysIO, SYS-36 Unbound

Environment, SYS-21 uncurry (CG), SYS-8 Unimplemented, LIB-1 union

BinarySet, LIB-6 IntSet, LIB-24 ORD SET, LIB-36 SplaySet, LIB-47

unique strings, LIB-33 uniqueSort

LIST SORT, LIB-27 ListMergeSort, LIB-28

unit, BASE-11 UNIX ENV, LIB-55 UNIX PATH, LIB-56 UnixEnv, LIB-55 UnixPath, LIB-56 unlink

SysIO, SYS-36 unroll (CG), SYS-7 UNSAFE, SYS-44 Unsafe, SYS-44 unzip

ListUtil, LIB-29 update

Array, BASE-3 Array2, LIB-2 ByteArray, BASE-6 DynamicArray, LIB-12 RealArray, BASE-26 Unsafe, SYS-44

updatef

Unsafe, SYS-44 use, GUIDE-2

Compile, SYS-15 use stream

Compile, SYS-15

VALspace, SYS-34 valSymbol, SYS-34 VECTOR, BASE-30 Vector, BASE-30 vector

Vector, BASE-30 version, LIB-1

System, SYS-40 versionName, LIB-1

WEAK, SYS-47

weak pointers, SYS-47

weak polymorphism, GUIDE-4 weakUnderscore

Control, SYS-18 whence

SysIO, SYS-36 wrap sysfn

CInterface, SYS-11 write

SysIO, SYS-36 writei

SysIO, SYS-36 writev

SysIO, SYS-36

xatoi StringCvt, LIB-49 xorb

Bits, BASE-4

Yacc, TOOL-13

Zip

ListUtil, LIB-29 zip

ListUtil, LIB-29